GO02

PI: Otterbein
Title: 0836+710

Session: Oct 1996 (session 3)
Frequency: 1.3 cm
Mode: B
Obs time (UT): 292 EU 16:30(18/10)-04:30(19/10)
Obs time (UT): 292 US 16:30(18/10)-04:30(19/10)
Stations: EbJbNtOn
Processor: Bonn
EU Mk3 Tapes/St.: 2

Comment


To give feedback on this experiment, click here

Experiment Feedback

Metsahovi (Kirsi Karlamaa, Ari Mujunen, Kaj Wiik at Mon Oct 21 02:58:48 EDT 1996) Minor failures
Antenna tracking computer problems 292 16:30--16:35. Weather was not particularly good, otherwise no known problems.

Effelsberg (T. Krichbaum at Tue Oct 22 10:59:00 EDT 1996) Success
The experiment should be correlated at Socorro. The corresponding statement in the EVN schedule is not correct. Barry Clark wants it to be correlated at AOC, because thin tapes were used and all VLBA sites participated.

Noto (Corrado Trigilio at Wed Oct 30 05:14:54 EST 1996) Success
No problems reported. Clear sky.

Onsala (A. Polatidis, J. Conway at Tue Nov 5 05:19:00 EST 1996) Success
No konwn problems. Cloudy throughout the experiment but no rain.

MPIfR (Walter Alef at Tue Sep 23 17:39:59 METDST 1997) Minor failures
Was correlated in Socorro!!!

\(Other\) ( at Thu Oct 17 01:08:06 CEST 2002) Minor failures